The lifetime of a certain type of battery is normally distributed with mean value 13 hours and standard deviation 1 hour. There are nine batteries in a package. What lifetime value is such that the total lifetime of all batteries in a package exceeds that value for only 5% of all packages? (Round your answer to two decimal places.) 
Answer #1

Let X be the random variable that denotes the lifetime of a certain type of battery.

X \sim Normal (\mu = 13, \sigma 2 = 1)

There are nine batteries in a package. Therefore, the sample size is n = 9.

According to the central limit theorem, if a random variable X is normally distributed with mean \mu and variance \sigma 2, then the distribution of the random variable \sum X is approximately normally distributed with mean n\mu and variance n\sigma2, where n is the sample size.

Using the central limit theorem, \sum X \sim Normal (n\mu = 9*13 = 117, n\sigma2 = 9*1 = 9)

Let the required lifetime value be c.

P(\sumX > c) = 0.05

\therefore P(Z > (c - 117) / \sqrt{9} ) = 0.05

\therefore 1 - P(Z < (c - 117) / \sqrt{9} ) = 0.05

\therefore P(Z < (c - 117) / \sqrt{9} ) = 0.95

From the standard normal table, P(Z < 1.645) = 0.95

\therefore (c - 117) / \sqrt{9} ) = 1.645

\therefore c - 117 = 3 * 1.645

\therefore c = 117 + 4.935

\therefore c = 121.935

Therefore, the required lifetime value is 121.94 hours.